A television set that originally cost $5500 was reduced in price by successive discounts of 15%, 30% and 20%. What was the equivalent single rate of discount?

Answer by ikleyn(52781)

.

The equivalent single rate of discount was  


    1 - (1-0.15)*(1-0.3)*(1-0.2) = 1 - 0.85*0.7*0.8 = 0.524,  or 52.4%.    ANSWER

------------

See the lesson
    - Percentage problems on chains of discounts
in this site.